x+3greather than or equal to 2x+1
Answer by checkley77(12844)

x+3>=2x+1
x-2x>=1-3
-x>=-2
x<=-2/-1 Note: when dividing by a negative value the > sign changes to a < sign.
x<=2 ans.
Proof:
Let x=1.9
1.9+3>=2*1.9+1
4.9>=3.8+1
4.9>=4.8